OpenStudy (anonymous):

Perform the division and write the answer without using negative exponents. -72r^2s^6t^3 over 27r^6s^3t^8

OpenStudy (anonymous):

it would be beneficial if you could factor these first

OpenStudy (anonymous):

I thought I answered this already? -72/27 = -8/3 r^2/r^6 = r^-4 or 1/r^4 and u go on like that...

OpenStudy (anonymous):

still lost

OpenStudy (anonymous):

can you factor these?

OpenStudy (anonymous):

kinda of

OpenStudy (anonymous):

well first factor them correctly, if you do this you will find terms that divide off.

OpenStudy (anonymous):

making the problem easier